The sum of the digits of a two digit number is 10.If 18 be subtracted from it the digits in the resulting number will be equal.Find the number.
Answer by vksarvepalli(154)

let the number be ab
we have a+b=10
and 10a+b-18 is a multiple of 11 (because it has the same digits)
now since a+b=10 we have few possibilities (91,82,73,64,19,28,37,46) now subtract 18 from these nos and see which no is divisible by 11
we will get the answer as 73
73-18=55